OpenStudy (blackops2luvr):

What is the length, in units, of the hypotenuse of a right triangle if each of the two legs is 2 units? 4 units 2 units 6sqrt units 8sqrt units @jim_thompson5910

OpenStudy (blackops2luvr):

Is it D?

jimthompson5910 (jim_thompson5910):

you already asked this

OpenStudy (blackops2luvr):

I know im checking my answer

jimthompson5910 (jim_thompson5910):

you need to solve for c a^2 + b^2 = c^2 2^2 + 2^2 = c^2

OpenStudy (blackops2luvr):

c is 8... 8^2

jimthompson5910 (jim_thompson5910):

8sqrt seems like a typo, do you mean sqrt(8) ?

OpenStudy (blackops2luvr):

Yes.

jimthompson5910 (jim_thompson5910):

yes, \[\Large c = \sqrt{8}\]

OpenStudy (blackops2luvr):

Okay/

OpenStudy (anonymous):

Hello I know your from OCVS and I would like to request your name. Your parents will be contacted and you will be removed from your course.

OpenStudy (blackops2luvr):

? i dont go to OCVS... i switched

OpenStudy (blackops2luvr):

This isnt even my account, im helping someone

OpenStudy (blackops2luvr):

Do you want me to let her know ??

OpenStudy (anonymous):

Lol its @Joel_the_boss I was sitting close to you on Tuesday. xDD

OpenStudy (anonymous):

Its cool man, I made this account to scare my friend. :P

OpenStudy (blackops2luvr):

XD WOOOOOOOOOW

OpenStudy (blackops2luvr):

I should do this, also @Joel_the_boss was one of ur fans so it was kinda obvious

OpenStudy (anonymous):

And know im going to scare everyone is OCVS XDDD You happened to be on so I targeted you first lol .

OpenStudy (blackops2luvr):

Brb, my name will be:

JoelTheBoss (joel_the_boss):

lol